OpenStudy (anonymous):

ue the function rule f(x) = n x |-1|^n find the output f(2) 4 3 -2 2

OpenStudy (anonymous):

@WhoaBaby14 @Well,Whatever @whpalmer4 @eSpeX @e.cociuba @Euler271 @EvonHowell @EsaaLoca @robtobey @RadEn

OpenStudy (anonymous):

this seems more like a f(n) than an f(x) f(2) = 2 * (-1)^2 = 2 * 1 = 2